# Scene: Procedural Controllable Large-scale Scene Generation via Large-language Models

### v1

@@ 4.2 Main Results
Furthermore, to compare with existing learning-based methods, we provide the qualitative comparison with several representing methods on 3D city generation, such as Persistent Nature [4], SceneDreamer [9],InfiniCity[31], and CityDreamer [52]. Fig. 5 shows the qualitative comparison results between the learning-based methods for city scene reconstruction work and Scene. From the results, we can observe that learning-based methods commonly suffer similar problems: low 3D consistency and building structural distortions. For example, PersistanNature [4] and InfiniCity [31] both appear to have severe deformation in the whole scene level. SceneDreamer [9] and CityDreamer [52] have better structure consistency, but the building quality is still relatively low. These factors limit their large-scale application in industry. In comparison, the proposed Scene generates highly realistic and well-structured urban scenes without the issues of structural distortions and layout defects compared to learning-based methods. These results demonstrate the effectiveness of our method for large-scale city scene generation.
@@
Aesthetic Evaluation. To better evaluate the generation quality of Scene, we collect the results of related works involving text-to-3D work and Blender-driven 3D generation. These results are subjected to aesthetic evaluation by a panel comprising of 35 voluntary contributors and 5 experts in 3D modeling, with the scoring criteria outlined in Section 4.1. As shown in Table 2, our scores for AS and AES surpass the second-highest scores by 2.10 and 0.9 points, respectively. Compared to the average level of other works, our project reaches a good level, indicating the high generation quality of Scene.
@@
To evaluate the consistency between text inputs and the generated assets, we calculate the CLIP similarity between input text and rendered images. To better illustrate the results, we utilize three different CLIP models for testing, including ViT-L/14 (V-L/14), ViT-B/16 (V-B/16) and ViT-B/32 (V-B/32), respectively. The detailed results are displayed in Table 2. We compare representative text-to-3D approaches (e.g. WonderJ [55], Text2Room [16], and DreamFusion [39]) and Blender-driven 3D generation works (e.g. BlenderGPT, 3D-GPT [45], and SceneCraft [19]). Although the similarity scores of text-to-3D methods are higher, it is reasonable because their training or optimization includes the text-to-image alignment process. Compared to the Blender-driven 3D generation works, Scene achieves the highest score, indicating its capability to accurately execute the input prompts and generate results.
@@
Efficiency Evaluation. To illustrate the efficiency of Scene, we provide the time required of our method compared with Infinigen [42] and human craft. The experiments are performed on a server equipped with dual Intel Xeon Processors (Skylake architecture), each with 20 cores, totaling 80 CPU cores. Additionally, we consult 3D PCG experts to determine the time needed to construct the same natural and urban scenes. The comparison results are shown in Table 3. From the results, we can observe that Scene is 7 times faster than Infinigen [42] in generating a scene with 150m 150m size. We also provide nearly 30 times faster than human experts creating a large-scale city by hand. This demonstrates the impressive efficiency of Scene in both large-scale natural scene and urban scene generation.
| 150m 150m | 500m 500m | 2.5km 2.5km | ||||
| Scene | City | Scene | City | Scene | City | |
| Human | 1h | 40min | 3h | 4h | - | >3w |
| Infinigen [42] | 14min | - | - | - | - | - |
| Scene (Ours) | 2min | 1.5min | 10min | 6min | - | 20h |
@@
4.3 Ablation Study
In order to analyze the impact of various components within the systematic template on the system, we conduct an ablation study based on the tree plugin in PCGBench. During the experiment, we utilize the dataset from Section 4.1 for testing, maintaining a consistent format. We incrementally add or remove different parts of the systematic template, using ER@1 and SR@1 as metrics to observe the impact of various components on the system. The results are shown in Table 4. It is evident from the results that augmenting the Task, Document, Examples and Role components contributes to an increase in ER@1 and SR@1. Among these, The inclusion of the Example component results in the most significant improvement, resulting in a maximum increase of 76.00% and 51.09% in ER@1 and SR@1 respectively. Conversely, the Role component has the least impact, with maximum increases of 2.00% and 2.59% in ER@1 and SR@1 after its addition. These results suggest that the setting of Task significantly impacts the performance of LLMs. By clarifying the goals and requirements of the task, the system can better understand the user’s intent and better meet the needs when generating output. Examples play an important role in designing proxy prompts. By providing concrete examples, the system can better understand the user’s intent and needs and produce high-quality output related to the input text. The inclusion of documentation is also important for designing proxy prompts. Documentation provides background information that can help the system better meet user expectations when generating output.
@@ 4.3 Ablation Study
4.4 Comparing with Different LLMs
@@
To investigate the performance of different variants of large language models (LLM) in Scene, we test public LLM APIs like gpt-3.5-turbo, gpt-4 and several external open-source LLMs in this subsection. To ensure the stability of LLM outputs, we set the temperature of LLM to 0 for all experiments. We conduct experiments on 3D scene and asset generation based on 50 scenario descriptions and 50 object descriptions in Sec. 4.1. The results are presented in Table 5. It is evident that gpt-4 delivers the most superior performance, with Mistral closely following as the second-best. Due to its superior performance and lower hardware requirements, the open-source Mistral is a highly appealing option. When compared to asset generation, the executability and success rates noticeably decline during the generation of large-scale natural scenes, a trend that can be attributed to the increased task complexity. In particular, as the number of components involved in the system expands, the LLM may face challenges in maintaining their accuracy. Nonetheless, our method exhibits consistent performance across different LLMs, maintaining high levels of executability and success rates.
| Model | Scene Generation | Asset Generation | ||
| ER@1 | SR@1 | ER@1 | SR@1 | |
| Llama2-7B[47] | 30.00 | 53.33 | 38.00 | 57.89 |
| Llama2-13B[47] | 44.00 | 59.09 | 54.00 | 66.66 |
| Mistral[24] | 76.00 | 68.42 | 94.00 | 85.11 |
| Gemma-2B[13] | 6.00 | 33.33 | 22.00 | 45.45 |
| Gemma-7B[13] | 36.00 | 55.56 | 68.00 | 73.53 |
| GPT-3.5-turbo[2] | 66.00 | 60.60 | 82.00 | 82.93 |
| GPT-4[37] | 86.00 | 86.05 | 96.00 | 85.42 |
5 Conclusion
@@
In this paper, we introduce an advanced framework for large-scale scene generation, denoted as Scene. This framework automatically generates high-quality procedural models based on designers’ textual descriptions. The proposed method consists of two main components: PCGBench and PCGPlanner Agent. PCGBench includes an extensive repository of accessible procedural assets and thousands of meticulously crafted API documents. On the other hand, the PCGPlanner Agent can generate executable actions for Blender to create controllable and precise 3D assets following user instructions. Utilizing Scene, we can generate a city covering 2.5 km 2.5 km with accurate layout and geometric structures. This significantly reduces the time required, from two weeks for several professional PCG engineers to just a few hours for an ordinary user. Our method’s capabilities in controllable large-scale scene generation and editing, including tasks like asset placement and season translation, are validated through extensive experiments.
Limitations Although the proposed Scene is effective for the LLM-driven procedural large-scale scene generation, it still faces several limitations: 1) The performance of our method relies on the pretrained LLM model. This reliance may constrain the framework from generalizing to a wider range of applications. 2) We have collected a limited number of assets and APIs, which limits the variety of generated scenes and degree of action space. In the future, it is promising to fine-tune the LLMs with professional modeling knowledge and empower them to be a professional modeling assistant. Collecting more resources for PCGBench can also enhance the scene fidelity and aesthetics.
6 Acknowledgement
@@
This work was supported in part by the National Key R&D Program of China (No. 2022ZD0116500), the National Natural Science Foundation of China (No. U21B2042, No. 62320106010), and in part by the 2035 Innovation Program of CAS and the InnoHK program.
Appendix 0.A Supplementary Materials
@@
0.A.1 PCGBench Collection
Blender Market offers a wide range of PCG plugins and 3D asset resources. To meet the demand for generating large-scale natural and urban scenes, we constructed PCGbench from four directions: Materials and Shading, Modeling, Rendering, and Asset Placement. Specifically, for generation-based PCG, we collected PCG models such as terrain, trees, flowers, rocks, and buildings. For material-based PCG, we gathered modifiable materials such as plant leaves, plant roots, terrain, buildings, snow, and sky. For asset placement-based PCG, we collected various placement PCG methods including slope-based, height-based, object-based, and geometric space-based placements. In terms of rendering-based PCG, we collected PCG resources related to rendering engines, rendering settings, and renderer optimizations. Additionally, for 3D asset resources, we collected a wide range of categories including buildings, vegetation, urban infrastructure, and textures sourced from the extensive realm of the internet.
Raw API collection: These PCG plugins are predominantly categorized into two types: one based on geometric node groups and the other based on Python code.It is necessary to gather API-related information from these PCG plugins, including API names, API function descriptions, API parameter descriptions, API parameter value ranges, and API parameter default values. For plugins based on geometric node groups, in order to obtain their original APIs, we need to test the geometric node groups of the plugins in Blender software and document their original APIs along with the corresponding information. For plugins based on Python code, we need to test the source code and record the functions within the plugin code as the original APIs, while also documenting the relevant information for these functions.
@@ 0.A.1 PCGBench Collection
PCGBench: In order to optimize the usage of PCG plugins in PCGPlanner, we need to simplify each plugin. The simplification rules are as follows:
@@
For generation-based plugins, we extract the APIs related to grid generation and remove irrelevant APIs. The criteria for successful extraction is the ability to independently generate the corresponding 3D grid using the plugin.
@@
For material-based plugins, we extract the APIs that provide material variations and remove irrelevant APIs. The criteria for successful extraction is the ability to independently change the corresponding material using the plugin.
@@
For placement-based plugins, we extract the APIs related to placement functionality and remove irrelevant APIs. Additionally, to accommodate multiple placement plugins, we have designed a universal collection format. The criteria for successful extraction is the ability to independently place objects in the current scene using the plugin.
@@
For rendering-based plugins, we extract the APIs that provide scene rendering capabilities and remove irrelevant APIs. The criteria for successful extraction is the ability to independently perform the rendering tasks for the entire 3D scene using the plugin.
Furthermore, to better integrate and utilize all the PCG plugins and 3D assets, we actively utilize Blender’s API to create some basic functions such as importing, deleting, moving, and deforming. These APIs can be accessed and called by PCGPlanner to achieve better results. 3D Asset Dataset: In this study, we have developed a comprehensive dataset consisting of a large collection of 3D assets, including 1908 meticulously crafted models and 1294 textures. To facilitate text-to-text and text-to-image retrieval for the CLIP model, we have extensively annotated each 3D asset and supplemented them with rendered images.
0.A.2 Algorithm
In this chapter, we provide the pseudocode for the entire framework implementation workflow to clearly illustrate the logical relationship between the loops between the various agents. First, we use a loop to check the termination condition, i.e. stop the execution of the algorithm when = True. In each loop, we first call the dispatch agent, pass the input and Prompt to the agent, and receive a set of rough plan objects . Next, we iterate over each object and use the specialist agent to further break it down into a detailed plan to get a set of . We then execute each detailed plan and call Agent_retrieval and CLIP to search for the corresponding assets and APIs in PCGBench. Finally, we evoke execution Agent to manipulate each retrieved API , until all the sub-task is completed and change the to True.
0.A.3 Examples of Prompt Details
@@
To provide a more intuitive demonstration of how agent prompts are designed and executed, we are presenting a series of examples. In Fig. 8 9 10 7, and 11, we show prompts for PCG plugins with different functionalities, along with the corresponding API calls for each PCG plugin. We also provide a tested example for each PCG component.
These images illustrate that our systematized prompt framework enables the successful execution of different tasks. Furthermore, the adoption of a unified prompt template for all agents simplifies maintenance and facilitates consistent invocation.
0.A.4 More Visualization Results

### v3

@@ Introduction
Generation Quality. We begin by showcasing several examples of our Scene for generating large-scale nature scenes and unbounded cities. The results are shown in Fig. 5. From the results, we can observe that the proposed Scene can produce highly realistic scenes in both natural scenes and cities. Moreover, the generated content is correctly corresponding to the provided texture descriptions. These demonstrate the power and effectiveness of our proposed LLM-driven automatic 3D scene generation framework. Fig. 6 shows the qualitative comparison results between the learning-based methods for city scene reconstruction work and Scene. From the results, we can observe that learning-based methods commonly suffer similar problems: low 3D consistency and building structural distortions. For example, PersistanNature (Chai et al. 2023) and InfiniCity (Lin et al. 2023b) both appear to have severe deformation in the whole scene level. SceneDreamer (Chen, Wang, and Liu 2023) and CityDreamer (Xie et al. 2024) have better structure consistency, but the building quality is still relatively low. These factors limit their large-scale application in industry. In comparison, the proposed Scene generates highly realistic and well-structured urban scenes without the issues of structural distortions and layout defects compared to learning-based methods. These results demonstrate the effectiveness of Scene for large-scale city scene generation.
| Method | AS | AES |
| Magic 3D (Lin et al. 2023a) | 4.48 | 3.50 |
| DreamFusion (Poole et al. 2022) | 4.55 | 3.60 |
| Text2Room (Höllein et al. 2023) | 5.73 | 6.10 |
| CityDreamer (Xie et al. 2024) | 5.47 | 6.80 |
| Infinigen (Raistrick et al. 2023) | 5.42 | 6.00 |
| 3D-GPT (Sun et al. 2023) | 4.94 | 6.20 |
| WonderJ (Yu et al. 2024) | 5.28 | 6.00 |
| Scene (Ours) | 7.83 | 7.70 |
@@
Aesthetic Evaluation. To better evaluate the generation quality of Scene, we collect the results of related works involving text-to-3D work and Blender-driven 3D generation. These results are subjected to aesthetic evaluation by a panel comprising of 35 voluntary contributors and 5 experts in 3D modeling, with the scoring criteria outlined in Section 4.1. As shown in Table 2, our scores for AS and AES surpass the second-highest scores by 2.10 and 0.9 points, respectively. Compared to the other works, our project reaches a good level, indicating the high generation quality of Scene.
@@
To evaluate the consistency between text inputs and the generated assets, we calculate the CLIP similarity between input text and rendered images. To better illustrate the results, we utilize three different CLIP models for testing, including ViT-L/14 (V-L/14), ViT-B/16 (V-B/16) and ViT-B/32 (V-B/32), respectively. The detailed results are displayed in Table 3. We compare representative text-to-3D approaches (e.g. WonderJ (Yu et al. 2024), Text2Room (Höllein et al. 2023), and DreamFusion (Poole et al. 2022)) and Blender-driven 3D generation works (e.g. BlenderGPT, 3D-GPT (Sun et al. 2023), and SceneCraft (Hu et al. 2024b)). Although the similarity scores of text-to-3D methods are higher, it is reasonable because their training or optimization includes the text-to-image alignment process. Compared to the Blender-driven 3D generation works, Scene achieves the highest score, indicating its capability to accurately execute the input prompts and generate results.
| V-L/14 | V-B/16 | V-B/32 | |
| WonderJ (Yu et al. 2024) | 18.78 | 25.70 | 25.45 |
| Text2Room (Höllein et al. 2023) | 23.51 | 30.10 | 29.29 |
| Magic 3D (Lin et al. 2023a) | 27.86 | 31.78 | 31.94 |
| DreamFusion (Poole et al. 2022) | 29.40 | 35.37 | 31.60 |
| BlenderGPT | 21.23 | 25.65 | 26.19 |
| 3D-GPT (Sun et al. 2023) | 18.67 | 25.80 | 25.59 |
| SceneCraft (Hu et al. 2024b) | 22.04 | 25.82 | 25.30 |
| Scene (Ours) | 22.82 | 27.82 | 26.89 |
| 150m 150m | 500m 500m | 2.5km 2.5km | ||||||
| Scene | City | Scene | City | Scene | City | |||
| Human | 1h | 40min | 3h | 4h | - | >3w | ||
| Infinigen | 14min | - | - | - | - | - | ||
| Scene | 2min | 1.5min | 10min | 6min | - | 20h | ||
@@
Efficiency Evaluation. To illustrate the efficiency of Scene, we provide the time required of our method compared with Infinigen (Raistrick et al. 2023) and human craft. The experiments are performed on a server equipped with dual Intel Xeon Processors (Skylake architecture), each with 20 cores, totaling 80 CPU cores. Additionally, we consult 3D PCG experts to determine the time needed to construct the same natural and urban scenes. The comparison results are shown in Table 4. From the results, we can observe that Scene is 7 times faster than Infinigen (Raistrick et al. 2023) in generating a scene with 150m 150m size. We also provide nearly 30 times faster than human experts creating a large-scale city by hand. This demonstrates the impressive efficiency of Scene in both large-scale natural scene and urban scene generation.
@@
Ablation Study
To analyze the impact of various components within the systematic template, we conduct an ablation study based on the tree plugin in PCGHub. During the experiment, we utilize the dataset from Section 4.1 for testing, maintaining a consistent format. We incrementally add or remove different parts of the systematic template, using ER@1 and SR@1 as metrics to observe the impact of various components on the system. The results are shown in Table 5. It is evident from the results that augmenting the Task, Document, Examples and Role components contributes to an increase in ER@1 and SR@1. Among these, the inclusion of the Example component results in the most significant improvement, resulting in a maximum increase of 76.00% and 51.09% in ER@1 and SR@1 respectively. Conversely, the Role component has the least impact, with maximum increases of 2.00% and 2.59% in ER@1 and SR@1 after its addition. These results suggest that the setting of Task significantly impacts the performance of LLMs. By clarifying the goals and requirements of the task, the system can better understand the user’s intent and better meet the needs when generating output. Examples play an important role in designing proxy prompts. By providing concrete examples, the system can better understand the user’s intent and needs and produce high-quality output related to the input text. Documentation provides background information that can help the system better meet user expectations when generating output.
Comparing with Different LLMs
@@
To investigate the performance of different variants of large language models (LLM) in Scene, we test public LLM APIs like gpt-3.5-turbo, gpt-4 and several external open-source LLMs in this subsection. To ensure the stability of LLM outputs, we set the temperature of LLM to 0 for all experiments. We conduct experiments on 3D scene and asset generation based on 50 scenario descriptions and 50 object descriptions in Sec. 4.1. The results are presented in Table 6. It is evident that gpt-4 delivers the superior performance, with Mistral closely following as the second-best. Due to its performance and lower hardware requirements, the open-source Mistral is a highly appealing option. When compared to asset generation, the executability and success rates noticeably decline during the generation of large-scale natural scenes, a trend that can be attributed to the increased task complexity. In particular, as the number of components involved in the system expands, the LLM may face challenges in maintaining their accuracy. Nonetheless, our method exhibits consistent performance across different LLMs, maintaining high levels of executability and success rates.
| Model | Scene Generation | Asset Generation | ||||
| ER@1 | SR@1 | ER@1 | SR@1 | |||
| Llama2-7B(Touvron et al. 2023) | 30.00 | 53.33 | 38.00 | 57.89 | ||
| Llama2-13B(Touvron et al. 2023) | 44.00 | 59.09 | 54.00 | 66.66 | ||
| Mistral(Jiang et al. 2023a) | 76.00 | 68.42 | 94.00 | 85.11 | ||
| Gemma-2B(Team et al. 2024) | 6.00 | 33.33 | 22.00 | 45.45 | ||
| Gemma-7B(Team et al. 2024) | 36.00 | 55.56 | 68.00 | 73.53 | ||
| GPT-3.5-turbo(Brown et al. 2020) | 66.00 | 60.60 | 82.00 | 82.93 | ||
| GPT-4(Achiam et al. 2023b) | 86.00 | 86.05 | 96.00 | 85.42 | ||
Conclusion
@@
In this paper, we introduce Scene, an advanced framework for large-scale scene generation that automatically creates high-quality procedural models from textual descriptions. The framework consists of two key components: PCGHub and PCGPlanner. PCGHub includes an extensive repository of accessible procedural assets and thousands of meticulously crafted API documents. Meanwhile, the PCGPlanner can generate executable actions for Blender to create controllable and precise 3D assets following user instructions. Scene can generate a 2.5 km × 2.5 km city with accurate layout and geometric structures, dramatically reducing production time from two weeks of professional work to just a few hours for an ordinary user. Our method demonstrates significant capabilities in controllable scene generation and editing, validated through extensive experiments.
Limitations. Although Scene effectively enables LLM-driven procedural large-scale scene generation, it has certain limitations: 1) Dependence on the pretrained LLM model may restrict its generalization across applications. 2) It is constrained by the inherent algorithms and rules of existing PCG modules. Expanding PCGHub’s resources could further enhance scene diversity and quality.
Acknowledgments
@@
This work was supported in part by the National Natural Science Foundation of China (No. U21B2042, No. 62320106010, No. 62072457), and the 2035 Innovation Program of CAS, and the InnoHK program.
Appendix A Supplementary Materials
@@
A. PCGHub Management
PCGHub offers an extensive collection of PCG modules and 3D asset resources designed to meet the demands of generating large-scale scenes. The management of PCGHub involves several key processes to ensure efficient operation and seamless integration of resources.
A.1.Encapsulation Standard
@@ A. PCGHub Management
To optimize the usage of PCG plugins, we need to simplify each plugin. The simplification rules are as follows:
@@
Generation Plugins: Extract APIs related to grid generation and remove irrelevant APIs, ensuring the ability to independently generate the corresponding 3D grid.
@@
Material Plugins: Extract APIs providing material variations and remove irrelevant APIs, ensuring the ability to independently modify the corresponding material.
@@
Placement Plugins: Extract APIs related to placement functionality and remove irrelevant APIs, ensuring the ability to place objects in the current scene.
@@
Rendering Plugins: Extract APIs providing scene rendering capabilities and remove irrelevant APIs, ensuring the ability to perform rendering tasks for the entire 3D scene.
A.2.PCG Encapsulation
PCG Encapsulation is a crucial process in PCGHub’s development, designed to integrate diverse PCG plugins—especially Blender node graph-based plugins—by establishing a standardized framework for their systematic utilization.
The encapsulation process meticulously examines each PCG plugin to extract its core program logic and input-output parameters. This approach ensures that each plugin’s essential functionality is preserved while enabling interoperability across different procedural generation tasks. The resulting encapsulated modules are stored in the blend file format, providing a standardized, reusable resource. Fig.8 illustrates the core program logic through a node graph visualization, showing the generation of tree roots and leaves, each with its corresponding node graph.
A.3.API Collection
PCG plugins are divided into two categories: those using geometric node groups and those based on Python code. To document these plugins, the following approach is taken: For plugins based on geometric node groups, we test these node groups in Blender to gather API details, including names, function descriptions, parameter descriptions, value ranges, and default values. For Python-based plugins, we analyze the source code to document functions, including names, purposes, parameters, ranges, default values, and constraints.
After defining the core logic and I/O parameters, the plugins are encapsulated as PCG modules in blend file format. Each module is then documented using Python code, covering the module’s name, description, inputs, input descriptions and types, outputs, and constraints. This documentation ensures each PCG module is effectively deployed for procedural content generation tasks. As shown in Fig. 11, the standardized Python template for documenting PCG modules streamlines and details each module, facilitating its integration into various generation workflows.
A.3.Rendering for Retrieval
The Rendering for Retrieval process automates the visualization of assets to facilitate their effective retrieval and integration into procedural content generation workflows. The procedure begins by computing the bounding box of a given 3D object, which is used to determine the object’s geometric center and form factor. These geometric parameters are essential for accurately positioning a tracking target within the scene. Specifically, the code generates an empty object at the asset’s geometric center, which serves as the focal point for subsequent camera configurations. Once configured, the camera is set as the active camera for rendering, and the system generates images that faithfully represent the asset’s key features. These images are then used for asset retrieval, enabling accurate matching with textual descriptions. As illustrated in Fig. 9, this process applies to building assets, generating rendered images that support accurate textual matching and retrieval within the system.
B. More Details
B.1. Algorithm
In this section, we present the pseudocode for the Scene framework implementation workflow, revealing the logical relationships across its stages. The process begins with a loop that checks the termination condition, halting execution when End_flag is True. Within each iteration, the workflow initiates by generating a set of rough plan objects from the Scene Decomposition stage, using input q and Prompt. Subsequently, each object is retrieved into the scene, followed by a RELATIONSHIP analysis to identify interactions among objects. Finally, asset placement is performed based on these relationships, continuing until all placement tasks are completed and End_flag is set to True.
B.2. Procedural Layout
PCGHub’s Procedural Layout Generators create diverse spatial configurations through five predefined layout types: Point Distribution, Grid-Based, Linear, Nested, and Region-Filling. These layouts model different real-world patterns, allowing for the simulation of complex environments. As illustrated in Fig. 10, the image demonstrate five examples of these layout types.
B.3. Examples of Prompt Details
@@
To provide a more intuitive demonstration of how prompts are designed and executed, we are presenting a series of examples. In Fig. 12, 13, 14, and 15, we show prompts for PCG plugins with different functionalities, along with the corresponding API calls for each PCG plugin. We also provide a tested example for each PCG component. These images illustrate that our systematized prompt framework enables the successful execution of different tasks. Furthermore, the adoption of a unified prompt template for all agents simplifies maintenance and facilitates consistent invocation.
Input: User query , knowledge documentation , The retrieved in the API collection, the retrieved in the asset collection , represents the Prompt of each Agent, represents the relationshop of each asset, .
Output: Modified API
B.4. More Visualization Results